OpenStudy (anonymous):

What is the equation of y = x3 with the given transformations? vertical compression by a factor of 1 over 7, horizontal shift 8 units to the left, reflection across the x-axis

OpenStudy (anonymous):

@ganeshie8 I know this is simple, but I'm so confused. Would you mind helping?

ganeshie8 (ganeshie8):

y = f(x) verticall compression by a factor of c : y = c*f(x) horizontal shift k units to the left : y = c*f(x+k) reflection across x axis : y = -c*f(x+k)

OpenStudy (triciaal):

@ganeshie8 just curious how do you write your quadratic in vertex form?

OpenStudy (anonymous):

Y=-1/7(x^3+8) @ganeshie8 ?

OpenStudy (anonymous):

@triciaal are you asking what vertex form is?

ganeshie8 (ganeshie8):

not quite

OpenStudy (triciaal):

no f(x) = a(x-h)^2 + k

ganeshie8 (ganeshie8):

y = x^3 verticall compression by a factor of 1/7 : y = (1/7)*f(x) horizontal shift 8 units to the left : y = (1/7)*(x+8)^3 reflection across x axis : y = -(1/7)*(x+8)^3

ganeshie8 (ganeshie8):

@triciaal quadratic ?

OpenStudy (anonymous):

Okay. So f(x) wasn't an example XD I understand now.. Thank you.

ganeshie8 (ganeshie8):

good. f(x) = x^3 f(x+k) = (x+k)^3

ganeshie8 (ganeshie8):

you just replace "x" by "x+k"

OpenStudy (triciaal):

@ganeshie8 oh because it's a cube function

OpenStudy (triciaal):

but the transformation results are the same

OpenStudy (anonymous):

Thank you, ganeshie.

OpenStudy (triciaal):

thanks